AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00 AR-00

Seventh Workshop on Automated Reasoning Bridging the Gap between

Theory and Practice

King’s College, London 20–21 July 2000

Following the highly successful series of Workshops on Automated Reasoning, this workshop will provide an informal forum for the automated reasoning community. This workshop series aims to bring together researchers from all areas of automated reasoning in order to foster links and facilitate cross-fertilisation of ideas among researchers from various disciplines; among researchers from academia, industry and government; and between theoreticians and practitioners.

ORGANISING COMMITTEE

• Robin D. Arthan, Lemma 1 rda@lemma-one.com

• Alan Bundy, Edinburgh A.Bundy@Edinburgh.ac.uk

• Tony Cohn, Leeds agc@scs.leeds.ac.uk

• Michael Fisher, Manchester M.Fisher@doc.mmu.ac.uk

• Alan M. Frisch, York (Chair of Org. Committee) frisch@cs.york.ac.uk

• Ian P. Gent, Glasgow ipg@cs.strath.ac.uk

• Andrew Ireland, Edinburgh air@cee.hw.ac.uk

• Manfred Kerber, Birmingham M.Kerber@cs.bham.ac.uk

• Hans Jürgen Ohlbach, London (Programme Chair) ohlbach@dcs.kcl.ac.uk

• Andrei Voronkov, Manchester voronkov@cs.man.ac.uk

• Toby Walsh, Glasgow tw@cs.strath.ac.uk TOPICS (not exclusive)

• Theorem proving in classical and non-classical logics

• Equational reasoning

• Unification

• Induction

• Logic programming

• Functional programming

• Constraint solving

• Formal methods for specifying, transforming and verifying systems, in particular safety-critical systems

• Non-monotonic reasoning

• Abduction

• Logic-based knowledge representation, in particular Description Logics.

Panel Sessions Proposals for panel sessions should be made as soon as possible. Please email the proposal to ohlbach@dcs.kcl.ac.uk.

Format of the Workshop There will be invited talks, panel sessions, short presentations of the papers, and a poster session. Each accepted paper will be introduced by a 5 minute overview talk. The technical details can then be presented in the poster session.

IMPORTANT DATES Receipt of submission (extended):  5 June 2000
Notification of authors:  3 July 2000
Receipt of camera-ready copy:  14 July 2000
Early bird registration deadline:  7 July 2000
WWW address  http://www.dcs.kcl.ac.uk/events/ARW/

SUBMISSION OF PAPERS

Two-Page Abstracts: We invite interested persons to submit a twopage abstract about recent work or work in progress, or a system description. Each submission should include the names and complete addresses (including email, when possible) of all authors. Correspondence will be sent to the first author, unless otherwise indicated. The main objective of the abstracts is to spread information about recent work in our community. Accepted abstracts will be published in informal workshop notes and be made available by WWW.

Full Papers In addition to the two-page abstracts, the participants are invited to submit full journal type papers. These papers will be treated as submissions to the Logic Journal of the IGPL (http://www.jigpal.oupjournals.org) They will be refereed and, if accepted, published by the journal as fast as possible.

Participation The workshop is intended to be an inclusive event, with participants encouraged from the broad spectrum covered by the field of automated reasoning. We encourage the participation of experienced researchers as well as those new to the field, especially students. A limited number of travel grants for students will be available. Students who want to apply for a travel grant should indicate this with their submission.

PROGRAMME CHAIR Hans Jürgen Ohlbach Dept. of Computer Science King’s College, Strand London WC2R 2LS, United Kingdom Tel.: +44 20 7848 2725 Fax: +44 20 7848 2851 Email: ohlbach@dcs.kcl.ac.uk CHAIR OF ORG. COMMITTEE Alan M. Frisch Department of Computer Science University of York York YO1 5DD United Kingdom Tel.: +44 1904 432745 Fax: +44 1904 432767 Email: frisch@cs.york.ac.uk LOCAL ARRANGEMENTS Ulrich Endriss (Tel: +44 20 7848 2985) Stefan Schlobach (Tel: +44 20 7848 2985) Odinaldo Rodrigues (Tel: +44 20 7848 2469) {endriss,schlobac,odinaldo}@dcs.kcl.ac.uk Dept. of Computer Science King’s College Strand London WC2R 2LS, Fax: +44 20 7848 2851 United Kingdom